# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Inhomogeneous Coefficients

### v1

@@ 3 Krylov-Based Adaptive-Rank Solver for Generalized Sylvester Equation
We propose an adaptive-rank algorithm for the generalized Sylvester equation (10). We assume a low-rank factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ at time $t^{(1)}=t^{(0)}+\Delta t$ also in the low-rank format. We represent this as:
@@
where $U_{0}\in\mathbb{R}^{N_{1}\times r_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r_{1}}$, and $V_{1}\in\mathbb{R}^{N_{2}\times r_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$ and $S_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$ are diagonal matrices with singular values at the corresponding times. Our method involves constructing the bases $U_{1}$ and $V_{1}$ and then projecting the original equation to determine the coefficients of $S_{1}$. A key aspect of the proposed adaptive-rank approach is that we do not form the full solutions explicitly. Instead, we seek bases in a dimension-by-dimension manner and then use them to perform Galerkin projection, resulting in a reduced system for the matrix of coefficients $S_{1}$. An SVD truncation procedure is then applied to update the orthonormal basis sets and realize an efficient low-rank representation of the updated solution. Key questions for solving the generalized Sylvester equation in the form of (10) that we address in this paper include:
  • How do we choose suitable subspaces to build $U_{1}$ and $V_{1}$ from $U_{0}$ and $V_{0}$?
@@
  • How can we check the residual of the generalized Sylvester equation without forming the matrices explicitly?
Our main innovation involves leveraging an approximate operator denoted by $\tilde{\mathscr{L}}$ (which is of the Sylvester form) to the original operator ${\mathscr{L}}$ in (9). $\tilde{\mathscr{L}}$ approximates the diagonal matrices $\Phi_{i}$ and $\Sigma_{i}$ representing pointwise multiplication of diffusion and advection coefficients, respectively, by scaled identities. This approximation not only allows us to construct dimension-wise Krylov bases $U_{1}$ and $V_{1}$ effectively and efficiently, but also enables effective preconditioning in the spirit of [25]. Then we project Equation (10) onto the Krylov bases, resulting in a reduced system for the coefficient matrix $S_{1}$. In this section, we introduce the main steps of our proposed algorithm as illustrated in Figure 1, highlighting the intuition behind our design and innovations along the way. Section 3.1 covers the construction of the approximated operator $\tilde{\mathscr{L}}$. Section 3.2 discusses the dimension-wise basis construction related to step 1 of the flowchart. In Section 3.3, we present the reduced equation for the matrix of coefficients $S_{1}$ and introduce our ACS preconditioner, corresponding to step 2 (i) in the flowchart. Lastly, Section 3.4 outlines the solution truncation mechanism and the residual computation, as described in steps 2(ii)-(iii) of the flowchart.
Figure 1: Flow-chart of the proposed implicit adaptive-rank algorithm for a generalized Sylvester equation.
@@ 3.1 Averaged-coefficient Approximate operator $\tilde{\mathscr{L}}$
Our approach incorporates ideas from [25] and [10] to construct an approximate operator $\tilde{\mathscr{L}}$ and efficiently solve the time-dependent inhomogeneous advection-diffusion equation using a projection-based adaptive-rank integration framework. The approximated operator $\tilde{\mathscr{L}}$ serves two purposes: (i) constructing the solution bases $U_{1}$ and $V_{1}$ as extended Krylov subspaces, and (ii) preconditioning the solution of the coefficient matrix $S_{1}$. We first construct a basis informed by both the original operator $\mathscr{L}$ and the approximate operator $\tilde{\mathscr{L}}$ (defined below). Secondly, we derive a corresponding projected (reduced) system that we solve iteratively using GMRES with a preconditioner derived from $\tilde{\mathscr{L}}$.
We seek an approximation to operator (9) that retains its spectral content and is easier to invert. As proposed in [25], the diagonal operators can be approximated by scaling with an identity matrix, where the scaling factor is the average of the diagonal entries. For instance, let $\Phi_{i}\approx{\alpha}_{i}I$ and $\Sigma_{i}\approx\gamma_{i}I$, where $\alpha_{i}$ and $\gamma_{i}$ are the average of the diagonal entries of $\Phi_{i}$ and $\Sigma_{i}$, respectively. This approximation yields the following mapping:
@@
A backward Euler temporal discretization using (14) yields the residual equation:
$\mathbf{R_{\tilde{\mathscr{L}}}}=\mathbf{F}_{1}-\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}=\mathbf{0},$
which leads to the Sylvester equation:
$(A_{1}+A_{3})\mathbf{F}_{1}+\mathbf{F}_{1}(A_{2}+A_{4})^{\top}=\mathbf{F}_{0}.$ (15)
Here:
@@
The solution to Equation (15) is an approximation to that of (10), but it can be efficiently computed using low-rank approximations [10, 33, 31, 30]. It can therefore be exploited as a viable preconditioner for a GMRES iterative solution of the generalized Sylvester equation (10). This was done in [25] with $\mathcal{O}(N^{3})$ computational complexity. In this work, we extend this approach by performing a Galerkin projection of both the original operator $\mathscr{L}$ and its approximation $\tilde{\mathscr{L}}$ on the low-rank bases before solving the generalized Sylvester equation with GMRES. This reduces the problem to a lower-dimensional subspace, where preconditioning can be applied more effectively on a smaller problem with $\mathcal{O}(r^{3})$ complexity, resulting in an overall computational complexity of $\mathcal{O}(Nr^{2}+r^{3})$.
@@ 3.2 Basis Construction
We begin by defining the generalized Krylov subspaces used for our basis construction at the $m$th iteration:
@@
$\kappa_{m}(M_{1},\cdots,M_{l},U)=[U,\underbrace{M_{1}U,\cdots,M_{l}U}_{\text{1st augmentation}},\ldots,\underbrace{M_{1}^{m+1}U,\cdots,M_{l}^{m+1}U}_{\text{mth augmentation}}].$ (18)
@@
where $M_{1},\cdots,M_{l}$ are operators of interest and their inverses arising in our generalized Sylvester equation, and $l$ is the number of operators used for each basis augmentation step.
@@
In our setting, the basis construction involves several operators from the original generalized Sylvester equation (10) and its approximation (15). We first consider the action of the diagonal operators, which have been averaged to obtain the approximated operators, $\Phi_{1}$, $\Sigma_{1}$, $\Phi_{2}$, and $\Sigma_{2}$. We have found that including these matrices is crucial for accelerating convergence. Using these operators, we construct a standard Krylov subspace. Additionally, we use the inverse operators from the approximated operators $A_{1:4}$ in (16)-(17) to form an inverted Krylov subspace. Finally, we construct dimension-wise extended Krylov subspaces for the approximated operator $\tilde{\mathscr{L}}$ using $P_{1}$ and $P_{2}$. The reasoning behind these choices of operators for subspace construction is as follows. Following [10], we construct an extended Krylov subspace in the $x$-direction for the approximated Sylvester equation (15) using $P_{1}$ and its inverse. To further enrich the basis, we include the inverses of the individual operators $A_{1}$ and $A_{3}$, as they may contain information not captured by the inverse of $P_{1}$. This carefully chosen set of operators results in a compact representation of the solution matrix, enriching the basis while keeping its dimension small.
Combining these, we define the extended Krylov subspaces in the $x$ direction as follows:
@@
$\kappa_{m}\left(P_{1},P_{1}^{-1},A_{1}^{-1},A_{3}^{-1},\Phi_{1},\Sigma_{1},U_{0}\right),$
We then apply an SVD-truncated QR, as outlined in Algorithm 1, to orthogonalize the subspace and obtain a basis $U_{1}^{(m)}\in\mathbb{{\bf R}}^{N_{1}\times r^{(m)}_{x}}$ and remove vectors that are closely linearly dependent. By leveraging the rank-revealing property of SVD, only those basis associated with significant singular values are retained. The same procedures are applied to the $y$-direction to obtain $V^{(m)}_{1}\in\mathbb{{\bf R}}^{N_{2}\times r^{(m)}_{y}}$. For notational simplicity, we will omit the superscript $(m)$ from now on, understanding that $U_{1}\in\mathbb{{\bf R}}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{{\bf R}}^{N_{2}\times r_{y}}$ refer to the bases of the Krylov subspaces at the $m$th iteration. This truncated orthonormalization strategy is particularly advantageous when the underlying processes (diffusion or advection) are weak. In such cases, the approximated coefficient matrices contribute minimally, reducing the impact of the operator application and enabling further basis truncation.
These constructed orthonormal bases will be used to represent the solution in the approximation subspace. The coefficients composing the linear combination of these basis vectors will be obtained by projecting the corresponding generalized Sylvester equation and evolving the matrix of coefficients $S_{1}$.
@@
2Compute the reduced QR decomposition: $\{Q,R\}=\texttt{QR}\left(\kappa_{m}\right)$;
3
4Perform reduced SVD decomposition: $T_{1}\tilde{R}T_{2}=\texttt{SVD}(R)$, where $\tilde{R}=\text{diag}(\sigma_{j})$;
@@
6Identify the last index $r$ such that $\sigma_{r}>\epsilon_{\kappa}$;
7
8Update the basis vectors: $Q\leftarrow QT_{1}(:,1:r)$;
9
Algorithm 1 SVD-truncated QR, $\mathscr{T}_{\epsilon_{\kappa}}$
@@ 3.3 Solving for the Matrix of Coefficients, $S_{1}$: ACS-preconditioned GMRES
With the orthonormal bases $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ at hand, we perform a Galerkin projection to derive a reduced generalized Sylvester equation for the matrix $S_{1}\in\mathbb{R}^{r_{x}\times r_{y}}$. The reduced Sylvester equation for $S_{1}$ arises from projecting the residual via the Galerkin condition $U_{1}^{\top}\mathbf{R}_{\mathscr{L}}V_{1}=0$, which yields:
@@
Solving Equation (19) using the Kronecker formulation $\mathcal{A}\text{vec}({{\bf S}_{1}})=\text{vec}({\tilde{{\bf B}}_{1}})$ with a direct solver requires a complexity of $\mathcal{O}(r^{6})$, where $r=\text{max}(r_{x},r_{y})$, which may be too expensive even for moderate $r$-values. Instead, we solve this equation using preconditioned GMRES, where the operator-times-vector operation is performed using a matrix-based strategy, with complexity $\mathcal{O}(r^{3})$. This is possible due to the Kronecker structure of our operator, yielding the equivalence between (19) and (20). Given a vector $s_{k}=\text{vec}(\mathbf{S}_{k})$, the product of the matrix $\mathcal{A}$ times a vector $s_{k}$ is organized as follows:
@@
The computational complexity of the GMRES solve also depends crucially on the total number of iterations needed to invert the linear system. To bound the number of GMRES iterations with respect to the basis dimension (rank) $r$, we require effective preconditioning, such that the eigenvalues of the preconditioned matrix are clustered and the number of iterations becomes independent of the system size $r$ [4]. Here, we propose a (left) preconditioner $\mathcal{P}$ such that the system:
@@
is well conditioned. The preconditioning step in GMRES requires applying $\mathcal{P}^{-1}$ to an arbitrary vector provided by the algorithm.
For fast GMRES convergence, $\mathcal{P}$ should approximate the left-hand side of Equation (20) well while being easier to invert than the original system. Here, we consider $\mathcal{P}=I_{r_{y}}\otimes\tilde{P}_{1}+\tilde{P}_{2}\otimes I_{r_{x}}$ arising from the Galerkin projection of Equation (15). In particular, let $U_{1}$ and $V_{1}$ be the orthonormal basis for the generalized Krylov subspaces, and let $\tilde{P}_{1}=U_{1}^{\top}P_{1}U_{1}$ and $\tilde{P}_{2}=V_{1}^{\top}P_{2}V_{1}$, then
$\mathcal{P}:\mathbf{S}\mapsto{\tilde{P}_{1}}\mathbf{S}+\mathbf{S}{\tilde{P}_{2}^{\top}}.$ (22)
To compute $\mathcal{P}^{-1}\hat{z}_{k}=z_{k}$, we solve the corresponding Sylvester equation:
@@
where $z_{k}=\text{vec}(\mathbf{Z}_{k})\in\mathbb{R}^{r_{x}r_{y}}$. Specific steps are as follows:
@@
    Solve $\mathcal{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\mathcal{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k}$ with complexity $\mathcal{O}(r^{3})$, for example with the Bartels Stewart algorithm [11].
@@
The proposed preconditioner effectively bounds the number of GMRES iterations required to solve Equation (19) independently of $r$, as will be shown later in this study, rendering the computational complexity of the $S$ matrix solve as $O(r^{3})$, much better than the expected $O(r^{6})$ scaling of a naïve implementation of GMRES.
3.4 Truncation, Efficient Residual Computation, and Algorithm Flowchart
At the end of each iteration, we need to compute the residual to accept or reject the low-rank approximation from Krylov subspaces. This process should take advantage of the low-rank representation of the solution while avoiding forming the residual on the full grid.
Once the solution ${\bf F}_{1}=U_{1}S_{1}V_{1}^{\top}$ is obtained, we apply truncation to control rank growth and reduce the residual computational load. Since $S_{1}$ from the GMRES solve is not necessarily diagonal, we follow the approach in [14] to convert it into a singular value form (also scaling as $\mathcal{O}(r^{3})$) and discard singular values below a threshold $\epsilon$. The truncation results in a solution ${\bf F}_{1}=U_{1}S_{1}V_{1}^{T}$, with $U_{1}\in\mathbb{R}^{N_{1}\times r_{1}}$, $V_{1}\in\mathbb{R}^{N_{2}\times r_{1}}$, and $S_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$, with $r_{1}\leq\text{min}(r_{x},r_{y})$. This procedure is summarized in Algorithm 2.
Following a similar procedure to [10], we devise a cost-effective way to compute the residual using the identity:

### v2

@@ 3 Krylov-Based Adaptive-Rank Solver for Generalized Sylvester Equation
We propose an adaptive-rank algorithm for the GSE (13). We assume a low-rank factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ at time $t^{(1)}=t^{(0)}+\Delta t$ also in the low-rank format, i.e.:
@@
where $U_{0}\in\mathbb{R}^{N_{1}\times r_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r_{1}}$, and $V_{1}\in\mathbb{R}^{N_{2}\times r_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$ and $S_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$ are diagonal matrices with singular values at the corresponding times. Projection methods involve constructing bases $U_{1}$ and $V_{1}$ and then projecting the original equation onto them to determine the coefficients of $S_{1}$. A key aspect of the proposed adaptive-rank approach is that we never form the full solutions explicitly. Instead, we seek bases in a dimension-by-dimension manner and then use them to perform a Galerkin projection. Key questions for solving the GSE that we address in this study include:
  • How do we choose suitable subspaces effectively to build $U_{1}$ and $V_{1}$ from $U_{0}$ and $V_{0}$?
@@
  • How can we check the residual of the GSE without forming the matrices explicitly?
Our main innovation leverages an approximate operator $\tilde{\mathscr{L}}$ (which is of the Sylvester form) to the original operator ${\mathscr{L}}$ in (12). This approximation not only allows us to construct dimension-wise Krylov bases $U_{1}$ and $V_{1}$ effectively and efficiently, but also enables effective preconditioning for solving the projected reduced system for $S_{1}$, in the spirit of [31]. In the following, we cover the construction of the approximated operator $\tilde{\mathscr{L}}$ in Section 3.1; we discuss the dimension-wise basis construction in Section 3.2; we present the reduced equation for the matrix of coefficients $S_{1}$ and introduce our ACS preconditioner in Section 3.3, and outline the residual computation in Section 3.4. Finally, Section 3.5 extends the algorithm to the case of multi-rank coefficients.
@@ 3.1 Averaged-Coefficient Approximate Operator $\tilde{\mathscr{L}}$
Our approach incorporates ideas from [31] and [14] to construct an approximate operator $\tilde{\mathscr{L}}$ and efficiently solve the time-dependent inhomogeneous advection-diffusion equation using a projection-based adaptive-rank integration framework. The approximated operator $\tilde{\mathscr{L}}$ serves two purposes: (i) constructing the solution bases $U_{1}$ and $V_{1}$ as xKrylov subspaces, and (ii) preconditioning the solution of the coefficient matrix $S_{1}$.
We seek an approximation to ${\mathscr{L}}$ that captures its spectral content and is easier to invert. As proposed in [31], the advection diagonal operators can be approximated by scaling an identity matrix, where the scaling factor is the average of the diagonal entries. In this work, we approximate not only the advection coefficients by a suitable constant multiple of the identity, but extend this strategy to the scalar diffusion coefficients by averaging them across the computational domain. Let $\Phi_{i}\approx{\alpha}_{i}I$ and $\Sigma_{i}\approx\gamma_{i}I$, where $\alpha_{i}$ and $\gamma_{i}$ are the averages of the diagonal entries of $\Phi_{i}$ and $\Sigma_{i}$, respectively. This approximation yields the following mapping:
@@
A backward Euler temporal discretization using (16) leads to the SE:
$(A_{1}+A_{3})\mathbf{F}_{1}+\mathbf{F}_{1}(A_{2}+A_{4})^{\top}=\mathbf{F}_{0},$ (17)
with:
@@
The solution to (17) is an approximation to that of (13), but it can be efficiently computed using low-rank approximations [14, 39, 37, 36]. It can therefore be exploited as a viable preconditioner for a GMRES iterative solution of the GSE (13). This was done in [31] with $\mathcal{O}(N^{3})$ computational complexity. In this work, we extend this approach by performing a Galerkin projection of both the original operator $\mathscr{L}$ and its approximation $\tilde{\mathscr{L}}$ on the low-rank bases before solving the GSE with GMRES. This reduces the problem to a lower-dimensional subspace, where preconditioning can be applied more effectively on a smaller $d$-dimensional problem with $\mathcal{O}(r^{d+1})$ complexity, resulting in an overall computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$.
@@ 3.2 Basis Construction
We begin by defining the xKrylov subspaces used for our basis construction at the $m$-th outer iteration as:
@@
$\kappa_{m}(M_{1},\cdots,M_{l},U)=\text{{\color[rgb]{0,0,0}Range}}\Big{\{}U,\underbrace{M_{1}U,\cdots,M_{l}U}_{\text{1st augmentation}},\ldots,\underbrace{M_{1}^{m+1}U,\cdots,M_{l}^{m+1}U}_{\text{mth augmentation}}\Big{\}}.$ (20)
@@
where $M_{1},\cdots,M_{l}$ are operators of interest and their inverses arising in our GSE, and $l$ is the number of operators used for each basis augmentation step.
@@
The xKrylov basis construction involves several operators from the original GSE (13) and its approximation (17). Following [14], we construct an xKrylov subspace in the $x$-direction for the approximated SE (17) using $P_{1}$ and its inverse. We also include the inverses of the individual operators $A_{1}$ and $A_{3}$, as they may contain information not captured by the inverse of $P_{1}$. Additionally, we consider the action of the diagonal operators $\Phi_{1}$, $\Sigma_{1}$, $\Phi_{2}$, and $\Sigma_{2}$, which are averaged to obtain $\tilde{\mathscr{L}}$. We have found that including all these matrices is effective in capturing the range of the solution. Using these operators, we construct a generalized Krylov subspace in the $x$-direction as follows:
@@
$\kappa_{m}\left(M_{1},\cdots,M_{6},U_{0}\right),\quad\text{where}$ $M_{1}=P_{1},$ $M_{2}=P_{1}^{-1},$ $M_{3}=A_{1}^{-1},$ $M_{4}=A_{3}^{-1},$ $M_{5}=\Phi_{1},$ $M_{6}=\Sigma_{1}.$
Depending on the structure of the GSE (2), $l$ may be large, and the generated subspace vectors at each augmentation may exhibit linear dependencies. To limit subspace inflation, we employ an SVD-based truncation strategy (Algorithm 1), such that only basis vectors corresponding to significant singular values are retained.
The construction of the dimension-wise xKrylov subspaces (20), following [39], is as follows. At each iteration $m$, Algorithm 2 augments the previously constructed basis $\mathcal{V}_{(m-1)}$ by applying powers of operators $M_{1}$ through $M_{l}$ to their corresponding basis blocks $V^{(1)},\dots,V^{(l)}$. Each block matrix $V^{(i)}$ may have a rank of at most $r_{0}$. We apply Algorithm 2 to orthogonalize the subspace and obtain a basis $\mathcal{U}_{(m)}\in\mathbb{R}^{N_{1}\times r^{(m)}_{x}}$ and $\mathcal{V}_{(m)}\in\mathbb{R}^{N_{2}\times r^{(m)}_{y}}$ in the $x$ and $y$ directions, respectively. For notational simplicity, we omit the calligraphic notation and subscript $(m)$ from now on, understanding that $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ refer to the bases of the Krylov subspaces at the $m$th iteration. The truncation step in line 2 of Algorithm 2 removes any linear dependencies within each block during the power iterations; line 2 applies a global truncation across all newly generated blocks, eliminating redundancies among them. As a result, the basis of the subspace is kept optimal in size.
@@
1 Compute the reduced QR decomposition: $\{Q,R\}=\texttt{QR}\left(\kappa_{m}\right)$;
2 Perform reduced SVD decomposition: $Z_{1}\tilde{R}Z_{2}=\texttt{SVD}(R)$, where $\tilde{R}=\text{diag}(\sigma_{j})$ Identify the last index $r$ such that $\sigma_{r}>\epsilon_{\kappa}$;
3 Update the basis vectors: $Q\leftarrow QZ_{1}(:,1:r)$;
Algorithm 1 SVD-truncated QR, $\mathscr{T}_{\epsilon_{\kappa}}$
Input: Operators $M_{1},\dots,M_{l}$; augmentation basis blocks $\{V^{(1)},\dots,V^{(l)}\}$; previous Krylov basis ${\mathcal{V}}_{(m-1)}$; tolerance $\epsilon_{\kappa}$.
Output: Updated Krylov basis $\mathcal{V}_{(m)}$; basis blocks $\{V^{(1)},\dots,V^{(l)}\}$ for the next iteration.
@@ 3 Update the basis vectors: $Q\leftarrow QZ_{1}(:,1:r)$;
2for $i=1,\dots,l$ do
3        $V^{(i)}\leftarrow M_{i}V^{(i)}$;
4        $\widetilde{V}^{(i)}\leftarrow\text{orthogonalize }V^{(i)}\text{ w.r.t. }V_{(m-1)}$ (e.g., modified Gram-Schmidt);
        $V^{(i)}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}(\widetilde{V}^{(i)})$ ;
         // truncate to $\epsilon_{\kappa}$ for next iteration using Algorithm 1
@@ 3        $V^{(i)}\leftarrow M_{i}V^{(i)}$;
[…]
$V^{\prime}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}(V^{\prime})$ ;
  // final orthogonalization and truncation using Algorithm 1
7 $\mathcal{V}_{(m)}\leftarrow[\mathcal{V}_{(m-1)},V^{\prime}]$;
Algorithm 2 Construction of xKrylov Subspace $\kappa_{m}^{\texttt{trunc}}$
We comment next on the performance of the xKrylov iteration. In principle, the number of xKrylov iterations may depend on the mesh size, $N$. For instance, [24, 4] theoretically show for the 2D and 3D Poisson equations that the dimension of the extended Krylov subspace (denoted by $r^{(m)}$ in the basis construction above) scales as $\mathcal{O}(\kappa^{1/4})\sim\mathcal{O}(N^{1/2})$, with $\kappa\sim N^{2}$ the condition number. While this provides an upper bound of the convergence rate (without considering solution smoothness), our numerical results suggest that it may be pessimistic in practice. In particular, for moderately tight tolerances, we find that the required Krylov subspace dimensions remain independent of grid resolution for a wide range of grid sizes $N$.
@@ 3.3 Solving for the Matrix of Coefficients, $S_{1}$: ACS-Preconditioned GMRES
With the orthonormal bases $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ at hand, we perform a Galerkin projection to derive a reduced GSE for the matrix $S_{1}\in\mathbb{R}^{r_{x}\times r_{y}}$. The reduced GSE for ${\bf S}_{1}$ arises from projecting the residual via the Galerkin condition $U_{1}^{\top}\mathbf{R}_{\mathscr{L}}V_{1}=0$, yielding:
@@
Solving the Kronecker formulation $\mathcal{A}\text{vec}({{\bf S}_{1}})=\text{vec}({\tilde{{\bf B}}_{1}})$ in (22) with a direct solver scales as $\mathcal{O}(r^{6})$ in 2D and $\mathcal{O}(r^{9})$ in 3D, where $r=\text{max}(r_{x},r_{y})$, which may be too expensive even for moderate $r$-values. Instead, we solve this equation iteratively using preconditioned GMRES. Instead of adapting GMRES to the matrix equation, we wrap GMRES around the Kronecker form, but exploiting the matrix-equation form (21) for both the $\mathcal{A}$-times-vector operation and the preconditioning step as indicated below, with overall complexity $\mathcal{O}(r^{3})$ in 2D (and $\mathcal{O}(r^{4})$ in 3D). For the GMRES operator-vector product, given a vector $s_{k}=\text{vec}(\mathbf{S}_{k})$, the product of the matrix $\mathcal{A}$ times a vector $s_{k}$ is performed as follows:
@@
For efficiency, GMRES requires preconditioning so that the eigenvalues of the preconditioned matrix are sufficiently clustered to deliver $r$-independent convergence rates  [7]. Here, we employ the ACS preconditioner, with $\mathcal{P}=I_{r_{y}}\otimes\tilde{P}_{1}+\tilde{P}_{2}\otimes I_{r_{x}}$ arising from the Galerkin projection of (17). The ACS preconditioner is expected to deliver optimal rank-independent, superlinear convergence (both observed in our numerical experiments), owing to its connection to the theory of compact-equivalent operators for diffusion-like equations, [^1] The averaged-coefficient advection-diffusion equation is compact-equivalent to the variable-coefficient one. which are known to be effective preconditioners for Krylov methods [1, 2].
We consider left preconditioning as:
@@
To compute $\mathcal{P}^{-1}\hat{z}_{k}=z_{k}$ for an arbitrary vector $\hat{z}_{k}$ during GMRES, we solve the corresponding SE:
@@
where $z_{k}=\text{vec}(\mathbf{Z}_{k})\in\mathbb{R}^{r_{x}r_{y}}$. Specific steps are:
@@
    Solve $\mathcal{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\mathcal{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k}$ with complexity $\mathcal{O}(r^{d+1})$, for example with the Bartels-Stewart algorithm [15].
@@
As a result, the $S$-matrix solve scales as $O(r^{d+1})$ overall, a great reduction in complexity compared to the $O(r^{3d})$ scaling of a naïve implementation of GMRES with the Kronecker form.
3.4 Truncation, Efficient Residual Computation
Once the solution ${\bf F}_{1}=U_{1}S_{1}V_{1}^{\top}$ is obtained, we apply truncation to control rank growth and reduce the residual computational load. Since $S_{1}$ resulting from the GMRES solve is not necessarily diagonal, we diagonalize it with SVD [19] [also scaling as $\mathcal{O}(r^{d+1})$] and discard singular values below a threshold $\epsilon$. The SVD truncation gives ${\bf F}_{1}=U_{1,\epsilon}S_{1,\epsilon}V_{1,\epsilon}^{\top}$, with $U_{1,\epsilon}\in\mathbb{R}^{N_{1}\times r_{1}}$, $V_{1,\epsilon}\in\mathbb{R}^{N_{2}\times r_{1}}$, and $S_{1,\epsilon}\in\mathbb{R}^{r_{1}\times r_{1}}$, with $r_{1}\leq\min(r_{x},r_{y})$ (Algorithm 3).
Following [14], we implement a cost-effective way to compute the residual using the identity: